# A Novel Image Similarity Metric for Scene Composition Structure

### v2

@@ Abstract
The rapid advancement of generative AI models necessitates novel methods for evaluating image quality that extend beyond human perception. A critical concern for these models is the preservation of an image’s underlying Scene Composition Structure (SCS), which defines the geometric relationships among objects and the background, their relative positions, sizes, orientations, etc. Maintaining SCS integrity is paramount for ensuring faithful and structurally accurate GenAI outputs. Traditional image similarity metrics often fall short in assessing SCS. Pixel-level approaches are overly sensitive to minor visual noise, while perception-based metrics prioritize human aesthetic appeal, neither adequately capturing structural fidelity. Furthermore, recent neural-network-based metrics introduce training overheads and potential generalization issues. We introduce the SCS Similarity Index Measure (SCSSIM), a novel, analytical, and training-free metric that quantifies SCS preservation by exploiting statistical measures derived from the Cuboidal hierarchical partitioning of images, robustly capturing non-object-based structural relationships. Our experiments demonstrate SCSSIM’s high invariance to non-compositional distortions, accurately reflecting unchanged SCS. Conversely, it shows a strong monotonic decrease for compositional distortions, precisely indicating when SCS has been altered. Compared to existing metrics, SCSSIM exhibits superior properties for structural evaluation, making it an invaluable tool for developing and evaluating generative models, ensuring the integrity of scene composition.
@@ 2 Scene Composition Structure (SCS)
In this section, we first formally define SCS in Section 2.1 and then discuss desirable properties of image similarity metrics for comparing SCS in Section $2.$
@@ 3.1 Cuboidal Partitioning
Let II be an image of width $w$ and height $h$. The sum of squared errors (SSE) $e$ of II is defined by
@@
where $p_{i}$ is the feature-vector of $i$-th pixel and $\mu=\frac{1}{|I|}\sum_{i\in I}p_{i}$ is the mean feature-vector of II. In this paper, the feature vectors are drawn from the RGB colour space. The CuPID algorithm recursively cuts II by finding the “strongest” horizontal or vertical lines with optimisation for minimising SSE.
Initially, the entire image II is considered as a partition, which has h1h-1 horizontal and w1w-1 vertical cuts in total as potential candidates for the “strongest” lines in the scene composition. Each cut splits the partition into two [sub] partitions $I_{1}$ and $I_{2}$ with SSE $e_{1}$ and $e_{2}$, respectively, defined by
@@
It can be mathematically proven that $e_{1}+e_{2}\leq e$. Thus, by splitting II into $I_{1}$ and $I_{2}$, the total SSE can only decrease. This decrease in the error value can be called the gain $g$ of the cut, defined by
@@
CuPID recursively continues this process, in a hierarchical fashion, with the [sub] partition of II, which can offer the maximum gain next, until a required NN number of final partitions are achieved with N1N-1 maximal cuts.
@@ 3.2 SCSSIM Similarity Metric
Now, we try to build a metric using the CuPID trees of two images being compared for SCS similarity. Naturally, the proposed similarity metric must conform to the desirable properties outlined in Section $2.$
@@
We build on this feature to design our metric. To compare an image II with a reference image $I_{0}$ we take the CuPID tree $T$ of $\tilde{N}$ cuts for $I_{0}$. When we apply the cuts of $T_{0}$ on $I_{0}$, we get the expected cumulative gain curve for $I_{0}$. If II is compositionally similar to $I_{0}$, it should produce a tree $T$ of $\tilde{N}$ cuts that can generate a similar curve when applied on $I_{0}$. Let the cumulative gain curves of applying $T_{0}$ and $T$ on $I_{0}$ be $c_{0}$ and $c$, respectively. When they are similarly structured, each term of $c$ should be close to those of $c_{0}$. So, their element-wise division should be fairly close to 1, i.e.,
@@
We expect $k_{i}$’s to have values very close to 0 for similar images. So, a good measure of their compositional distance would be their mean $\bar{k}=\frac{1}{\tilde{N}}\sum_{i=1}^{\tilde{N}}k_{i}$. However, as $\bar{k}$ is not necessarily bounded, we apply exponential decay on it. Thus, the metric $\mathrm{M_{fr}}(I,I_{0})$, which tells us how similar II is to $I_{0}$ is defined by
@@
where $\lambda$ is the decay rate. A quadratic decay has been used to avoid negative $\bar{k}$ values and to penalize large differences. Through a sensitivity study for different $\lambda$’s, which is not reported in this paper for page restrictions, $\lambda=25$ is empirically found to give good results, which has been used in the experiments reported in Section 4.
@@ 4 Experimental Results
We highlight the result for SCSSIM in Table 1 on the images in Fig. 2. Unlike other similarity metrics, SCSSIM values are consistent with the expectations made in Section $2.$ We have empirically found $\tilde{N}=64$ to be the most promising. Fewer cuts tend to misrepresent the curved or slanted composition structures, and more cuts start to capture textural details in the image that are not part of SCS. All further results are provided for CuPID trees for this $\tilde{N}$.

### v3

@@ Abstract
The rapid advancement of generative AI models necessitates novel methods for evaluating image quality that extend beyond human perception. A critical concern for these models is the preservation of an image’s underlying Scene Composition Structure (SCS), which defines the geometric relationships among objects and the background, their relative positions, sizes, orientations, etc. Maintaining SCS integrity is paramount for ensuring faithful and structurally accurate GenAI outputs. Traditional image similarity metrics often fall short in assessing SCS. Pixel-level approaches are overly sensitive to minor visual noise, while perception-based metrics prioritize human aesthetic appeal, neither adequately capturing structural fidelity. Furthermore, recent neural-network-based metrics introduce training overheads and potential generalization issues. We introduce the SCS Similarity Index Measure (SCSSIM), a novel, analytical, and training-free metric that quantifies SCS preservation by exploiting statistical measures derived from the Cuboidal hierarchical partitioning of images, robustly capturing non-object-based structural relationships. Our experiments demonstrate SCSSIM’s high invariance to non-compositional distortions, accurately reflecting unchanged SCS. Conversely, it shows a strong monotonic decrease for compositional distortions, precisely indicating when SCS has been altered. Compared to existing metrics, SCSSIM exhibits superior properties for structural evaluation, making it an invaluable tool for developing and evaluating generative models, ensuring the integrity of scene composition. See code.
@@ 2 Scene Composition Structure (SCS)
In this section, we first formally define SCS in Section 2.1 and then discuss desirable properties of image similarity metrics for comparing SCS in Section 2.2.
@@ 3.1 Cuboidal Partitioning
Let $I$ be an image of width $w$ and height $h$. The sum of squared errors (SSE) $e$ of $I$ is defined by
@@
where $p_{i}$ is the feature-vector of $i$-th pixel and $\mu=\frac{1}{|I|}\sum_{i\in I}p_{i}$ is the mean feature-vector of $I$. In this paper, the feature vectors are drawn from the RGB colour space. The CuPID algorithm recursively cuts $I$ by finding the “strongest” horizontal or vertical lines with optimisation for minimising SSE.
Initially, the entire image $I$ is considered as a partition, which has h1h-1 horizontal and w1w-1 vertical cuts in total as potential candidates for the “strongest” lines in the scene composition. Each cut splits the partition into two [sub] partitions $I_{1}$ and $I_{2}$ with SSE $e_{1}$ and $e_{2}$, respectively, defined by
@@
It can be mathematically proven that $e_{1}+e_{2}\leq e$. Thus, by splitting $I$ into $I_{1}$ and $I_{2}$, the total SSE can only decrease. This decrease in the error value can be called the gain $g$ of the cut, defined by
@@
CuPID recursively continues this process, in a hierarchical fashion, with the [sub] partition of $I$, which can offer the maximum gain next, until a required $N$ number of final partitions are achieved with N1N-1 maximal cuts.
@@ 3.2 SCSSIM Similarity Metric
Now, we try to build a metric using the CuPID trees of two images being compared for SCS similarity. Naturally, the proposed similarity metric must conform to the desirable properties outlined in Section 2.2.
@@
We build on this feature to design our metric. To compare an image $I$ with a reference image $I_{0}$ we take the CuPID tree $T$ of $\tilde{N}$ cuts for $I_{0}$. When we apply the cuts of $T_{0}$ on $I_{0}$, we get the expected cumulative gain curve for $I_{0}$. If $I$ is compositionally similar to $I_{0}$, it should produce a tree $T$ of $\tilde{N}$ cuts that can generate a similar curve when applied on $I_{0}$. Let the cumulative gain curves of applying $T_{0}$ and $T$ on $I_{0}$ be $c_{0}$ and $c$, respectively. When they are similarly structured, each term of $c$ should be close to those of $c_{0}$. So, their element-wise division should be fairly close to 1, i.e.,
@@
We expect $k_{i}$’s to have values very close to 0 for similar images. So, a good measure of their compositional distance would be their mean $\bar{k}=\frac{1}{\tilde{N}}\sum_{i=1}^{\tilde{N}}k_{i}$. However, as $\bar{k}$ is not necessarily bounded, we apply exponential decay on it. Thus, the metric $\mathrm{M_{fr}}(I,I_{0})$, which tells us how similar $I$ is to $I_{0}$ is defined by
@@
where $\lambda$ is the decay rate. A quadratic decay has been used to avoid negative $\bar{k}$ values and to penalize large differences. Through a sensitivity study for different $\lambda$’s, which is not reported in this paper for page restrictions, $\lambda=5$ is empirically found to give good results, which has been used in the experiments reported in Section 4.
@@ 4 Experimental Results
We highlight the result for SCSSIM in Table 1 on the images in Fig. 2. Unlike other similarity metrics, SCSSIM values are consistent with the expectations made in Section 2.2. We have empirically found $\tilde{N}=64$ to be the most promising. Fewer cuts tend to misrepresent the curved or slanted composition structures, and more cuts start to capture textural details in the image that are not part of SCS. All further results are provided for CuPID trees for this $\tilde{N}$.